) Given the equation 2x – 6y = 12:
a) Rewrite the equation in slope-intercept form.



b) Write an equation in point-slope form of the line through the point (-4, 1) parallel to the line with the equation given above.



c) Convert the equation found in part (b) to slope-intercept form.



d) Convert the equation found in part (c) to standard form, Ax + By = C, where A, B, and C are integers.

Answer by Alan3354(69443)

Given the equation 2x – 6y = 12:
a) Rewrite the equation in slope-intercept form.
Solve for y.
=============
b) Write an equation in point-slope form of the line through the point (-4,1) parallel to the line with the equation given above.
Use y-y1 = m*(x-x1) where m = slope and (x1,y1) is the point.
-------------
c) Convert the equation found in part (b) to slope-intercept form.
Do that.
-----------
d) Convert the equation found in part (c) to standard form, Ax + By = C, where A, B, and C are integers.
Do that.